Kabz is 12 years older than Jane. In 10 years twice Kabz's age decrease by Jane's will be 35. Find their ages now.: Kabz is 5 years older than Jane. In 10 years twice Kabz's age decrease by Jane's will be 35. Find their ages now.

Answer by ankor@dixie-net.com(12692)

Write an expression for each phrase
:
"Kabz is 12 years older than Jane."
K = J + 12
:
"In 10 years":
K+10
and
J+10
:
"In 10 yrs twice Kabz's age decrease by Jane's will be 35.
2(K+10) - (J+10) = 35
2K + 20 - J - 10 = 35
2K +10 - J = 35
2K + 10 - 35 = +J
Therefore:
J = 2K - 25
:
Find their ages now.:
Substitute (J+12) for K
J = 2(J+12) - 25
J = 2J + 24 - 25
J - 2J = -1
-J = -1
J = 1 yr old now
:
That makes K = 1+12 = 13 yrs old now
:
Check solution using "In 10 yrs twice Kabz's age decrease by Jane's will be 35.
2(23) - 11 = 35
:
:
Kabz is 5 years older than Jane. In 10 years twice Kabz's age decrease by Jane's will be 35. Find their ages now.
:
K = J + 5
Take the equation we created above:
J = 2K - 25
:
Now substitute (J+5) for K in the above equation:
J = 2(J+5) - 25
J = 2J + 10 - 25
J = 2J - 15
J - 2J = -15
-J = -15
J = 15 yrs old now
Therefore:
K = 15 + 5 = 20 yrs old now
:
Check solution using: "In 10 yrs twice Kabz's age decrease by Jane's will be 35"
2(30) - 25 = 35
:
:
Did this help?


: